How Long will it take to walk 17 miles (27 km)?
Answer by jim_thompson5910(35256)

It will depend on the speed.

If you are walking at a speed of 1 mph, then it will take 17/1 = 17 hours

If you are walking at a speed of 2 mph, then it will take 17/2 = 8.5 hours

If you are walking at a speed of 3 mph, then it will take 17/3 = 5.667 hours

If you are walking at a speed of 4 mph, then it will take 17/4 = 4.25 hours

etc etc